OpenStudy (anonymous):

HELP! if secx=-5 compute : tan(-t)+cot(-t) please work it out i am lost.

OpenStudy (anonymous):

i dont understand the question - should it be tan(-x) + cot(-x)?

OpenStudy (anonymous):

thats wa i thought as well, but that is exactly what the paper has written down

OpenStudy (anonymous):

Use Sec^2 = 1 + Tan^2 (must be a mistake on paper, either all x's or all t's).

OpenStudy (anonymous):

ok gotcha

OpenStudy (anonymous):

did you solve this one ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

sec(x)=-5 --> cos(x)=-0.2 Some trig identities: cos(-x)=cos(x) sin(-x)=-sin(x) (sin x)^2+(cos x)^2=1 --> (sin x)^2 = 1 - (cos x)^2 = 1 - (-0.2)^2 --> sin x = sqrt( 1 - (-0.2)^2 ) = sqrt(24/25) = 0.2*sqrt(6) So now we have: sin(x) = 0.2 * sqrt(6) cos(x) = -0.2 Next, tan(-x) + cot(-x) = sin(-x)/cos(-x) + cos(-x)/sin(-x) = -sin(x)/cos(x) - cos(x)/sin(x) = -( sin(x)/cos(x) + cos(x)/sin(x)) = -( (0.2*2*sqrt(6))/(-0.2) + (-0.2)/(0.2*2*sqrt(6))) = -( -2*sqrt(6) - 1/(2*sqrt(6)) = 2*sqrt(6) + 1/(2*sqrt(6)) = (24 + 1)/(2*sqrt(6)) = 25 / (2*sqrt(6)) = 25 * sqrt(6) / 12 So: tan(-x)+cot(-x) = 25*sqrt(6)/12

OpenStudy (anonymous):

Making a lot of work for yourself, u could use my suggestion in first post: -tanx -cotx = -tan x -1/tanx = (-tansx - 1)/tanx = -(1+tan^2x)/tan x = sec^2x/tanx = 25/tan x and use the same formula again to get 25/sqrt24 which is same answer as u.

OpenStudy (anonymous):

estudier - which formula did use again to get from 25/tan x to 25/sqrt24 ?

OpenStudy (anonymous):

Same one Use Sec^2 = 1 + Tan^2 (take sqrt)

OpenStudy (anonymous):

yep like it better than mine .. :)

OpenStudy (anonymous):

gotcha yea im taking precal in the summer with a horrible professor (larry lace, moorpark college check em out rate my profe.) so basically im teaching myself so that's y im a few steps behind.

OpenStudy (anonymous):

A good resource for self-learning: http://www.khanacademy.org/ e.g.: search for "Trigonometry" , "Precalculus" on that page. Lot's of good videos with friendly and clear explanations. There are also practice questions you can do (click "PRACTICE" on the top of the main page), where the system assesses your level and drills you further in areas you need improve etc..